xx 1

Marcus Ansorg <marcus.ansorg@aei.mpg.de> Erik Schnetter <schnetter@cct.lsu.edu>

May 18 2004

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11 admxtol 10 Scope: restricted01 REAL

11 Description: Tolerance of ADM masses when givexbarexmass=no

11 Range 10 01 Default: 1.0e-10
11 (0:*)  10

11

11 centerxoffset  10 Scope: restricted01  REAL

11 Description: offset b=0 to position (x,y,z)

11 Range 10 01 Default: 0.0
11 (*:*)  10

11

11 doxinitialxdebugxoutput10 Scope: restricted01  BOOLEAN

11 Description: Output debug information about initial guess

11                      10                01  Default: no

11

11 doxresiduumxdebugxoutput 10 Scope: restricted01  BOOLEAN

11 Description: Output debug information about the residuum

11                      10                01  Default: no

11

11 givexbarexmass  10 Scope: restricted01  BOOLEAN

11 Description: User provides bare masses rather than target ADM masses

11                      10                01  Default: yes

11

11 gridxsetupxmethod  10 Scope: restricted01  KEYWORD

11 Description: How to fill the 3D grid from the spectral grid

11 Range 10 01 Default: Taylor expansion
11 Taylor expansion  10 use a Taylor expansion about the nearest collocation point (fast, but might be inaccurate)
11 evaluation  10 evaluate using all spectral coefficients (slow)

11

11 initialxlapsexpsixexponent 10 Scope: restricted01  REAL

11 Description: Exponent n for psiˆ- n initial lapse profile

11 Range 10 01 Default: -2.0
11 (*:*)  10 Should be negative

11

11 keepxuxaround  10 Scope: restricted01  BOOLEAN

11 Description: Keep the variable u around after solving

11                      10                01  Default: no

11

11 multiplyxoldxlapse  10 Scope: restricted01  BOOLEAN

11 Description: Multiply the old lapse with the new one

11                      10                01  Default: no

11

11 newtonxmaxit  10 Scope: restricted01  INT

11 Description: Maximum number of Newton iterations

11  Range  10                01  Default: 5

11 0:*  10

11

11 newtonxtol  10 Scope: restricted01  REAL

11 Description: Tolerance for Newton solver

11 Range 10 01 Default: 1.0e-10
11 (0:*)  10

11 npointsxa  10 Scope: restricted01  INT

11 Description: Number of coefficients in the compactified radial direction

11  Range  10                01  Default: 30

11 4:*  10

11

11 npointsxb  10 Scope: restricted01  INT

11 Description: Number of coefficients in the angular direction

11  Range  10                01  Default: 30

11 4:*  10

11

11 npointsxphi  10 Scope: restricted01  INT

11 Description: Number of coefficients in the phi direction

11 Range 10 01 Default: 16
11 4:*:2  10

11

11 parxb  10 Scope: restricted01  REAL

11 Description: x coordinate of the m+ puncture

11 Range 10 01 Default: 1.0
11 (0.0:*)  10

11

11 parxmxminus  10 Scope: restricted01  REAL

11 Description: mass of the m- puncture

11 Range 10 01 Default: 1.0
11 0.0:*)  10

11

11 parxmxplus  10 Scope: restricted01  REAL

11 Description: mass of the m+ puncture

11 Range 10 01 Default: 1.0
11 0.0:*)  10

11

11 parxpxminus  10 Scope: restricted01  REAL

11 Description: momentum of the m- puncture

11 Range 10 01 Default: 0.0
11 (*:*)  10

11

11 parxpxplus  10 Scope: restricted01  REAL

11 Description: momentum of the m+ puncture

11 Range 10 01 Default: 0.0
11 (*:*)  10

11

11 parxsxminus  10 Scope: restricted01  REAL

11 Description: spin of the m- puncture

11 Range 10 01 Default: 0.0
11 (*:*)  10

11

11 parxsxplus  10 Scope: restricted01  REAL

11 Description: spin of the m+ puncture

11 Range 10 01 Default: 0.0
11 (*:*)  10

11

11 rescalexsources  10 Scope: restricted01  BOOLEAN

11 Description: If sources are used - rescale them after solving?

11                      10                01  Default: yes

11 solvexmomentumxconstraint 10 Scope: restricted01  BOOLEAN

11 Description: Solve for momentum constraint?

11                      10                01  Default: no

11

11 swapxxz  10 Scope: restricted01  BOOLEAN

11 Description: Swap x and z coordinates when interpolating, so that the black holes are separated in the z direction

11                      10                01  Default: no

11

11 targetxmxminus  10 Scope: restricted01  REAL

11 Description: target ADM mass for m-

11 Range 10 01 Default: 0.5
11 0.0:*)  10

11

11 targetxmxplus  10 Scope: restricted01  REAL

11 Description: target ADM mass for m+

11 Range 10 01 Default: 0.5
11 0.0:*)  10

11

11 tpxepsilon  10 Scope: restricted01  REAL

11 Description: A small number to smooth out singularities at the puncture locations

11  Range  10                01  Default: 0.0

11 0:*  10

11

11 tpxextendxradius  10 Scope: restricted01  REAL

11 Description: Radius of an extended spacetime instead of the puncture

11 Range 10 01 Default: 0.0
11 0:*  10 anything positive, should be smaller than the horizon

11

11 tpxtiny  10 Scope: restricted01  REAL

11 Description: Tiny number to avoid nans near or at the pucture locations

11 Range 10 01 Default: 0.0
11 0:*  10 anything positive, usually very small

11

11 usexexternalxinitialxguess 10 Scope: restricted01  BOOLEAN

11 Description: Set initial guess by external function?

11                      10                01  Default: no

11

11 usexsources  10 Scope: restricted01  BOOLEAN

11 Description: Use sources?

11                      10                01  Default: no

11

11 verbose  10 Scope: restricted01  BOOLEAN

11 Description: Print screen output while solving

11                      10                01  Default: no

11

11 outxdir  10 Scope: shared from IO01STRING

7 Interfaces

General Implements: twopuncturesx

Inherits: admbasex coordinatesx

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 puncturexu 10 puncturexu 01 compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  energy  10                    01  compact 11 0
11                    10  E  01 description 11 ADM energy of the Bowen-York spacetime
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 angularxmomentum 10                    01  compact 11 0
11                    10  J1  01 description 11 Angular momentum of the Bowen-York spacetime
11                    10  J2  01 dimensions 11 0
11                    10  J3  01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  barexmass  10                    01  compact 11 0
11                    10  mp  01 description 11 Bare masses of the punctures
11                    10  mm  01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 puncturexadmxmass 10                    01  compact 11 0
11                    10  mpxadm  01 description 11 ADM masses of the punctures (measured at the other spatial infinities)
11                    10  mmxadm  01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn SpacetimeX/TwoPuncturesX. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 x 11 Conditional:
11 x 11 energy angularxmomentum puncturexadmxmass barexmass
11 x 11 puncturexu

11 x 11 x

Scheduled Functions

CCTKxPARAMCHECK (conditional) x twopuncturesxxparamcheck x check parameters and thorn needs x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

ADMBaseXxInitialData (conditional) x twopuncturesxxgroup x twopuncturesx initial data group x

11 x 11 Type: 11 group

11

TwoPuncturesXxGroup (conditional) x twopuncturesxxtwopunctures x create puncture black hole initial data x

11 x 11 Language: 11 c
11 x 11 Reads:  11 coordinatesx::vertexxcoords(everywhere)
11 x 11 Storage: 11 puncturexu
11 x 11 Type:  11 function
11 x 11 Writes: 11 admbasex::metric(everywhere)
11 x 11 x  11 admbasex::curv(everywhere)
11 x 11 x  11 admbasex::lapse(everywhere)
11 x 11 x  11 puncturexu(everywhere)

11 x 11 x  11 mp
11 x 11 x  11 mm

11 x 11 x  11 mpxadm
11 x 11 x  11 mmxadm

11 x 11 x  11 e
11 x 11 x  11 j1
11 x 11 x  11 j2
11 x 11 x  11 j3

11

TwoPuncturesXxGroup (conditional) x twopuncturesxxmetadata x output twopuncturesx metadata x

11 x 11 After:  11 twopuncturesx
11 x 11 Language: 11 c
11 x 11 Options: 11 global
11 x 11 Reads:  11 mp

11 x 11 x  11 mm

11 x 11 x  11 mpxadm
11 x 11 x  11 mmxadm

11 x 11 x  11 e
11 x 11 x  11 j1
11 x 11 x  11 j2
11 x 11 x  11 j3

11 x 11 Type:  11 function